# Index Terms:

### v1

@@ Abstract
Edge intelligence (EI) allows resource-constrained edge devices (EDs) to offload computation-intensive AI tasks (e.g., visual object detection) to edge servers (ESs) for fast execution. However, transmitting high-volume raw task data (e.g., 4K video) over bandwidth-limited wireless networks incurs significant latency. While EDs can reduce transmission latency by degrading data before transmission (e.g., reducing resolution from 4K to 720p or 480p), it often deteriorates inference accuracy, creating a critical accuracy-latency tradeoff. The difficulty in balancing this tradeoff stems from the absence of closed-form models capturing content-dependent accuracy-latency relationships. Besides, under bandwidth sharing constraints, the discrete degradation decisions among the EDs demonstrate inherent combinatorial complexity. Mathematically, it requires solving a challenging black-box mixed-integer nonlinear programming (MINLP). To address this problem, we propose LAB, a novel learning framework that seamlessly integrates deep reinforcement learning (DRL) and Bayesian optimization (BO). Specifically, LAB employs: (a) a DNN-based actor that maps input system state to degradation actions, directly addressing the combinatorial complexity of the MINLP; and (b) a BO-based critic with an explicit model built from fitting a Gaussian process surrogate with historical observations, enabling model-based evaluation of degradation actions. For each selected action, optimal bandwidth allocation is then efficiently derived via convex optimization. Numerical evaluations on real-world self-driving datasets demonstrate that LAB achieves near-optimal accuracy-latency tradeoff, exhibiting only 1.22% accuracy degradation and 0.07s added latency compared to exhaustive search. Notably, it outperforms conventional DRL with 3.29% higher accuracy and 42.60% lower latency, demonstrating its advantageous performance in handling black-box optimization problems.
@@ I Introduction
Edge inference (EI) enables resource-constrained edge devices (EDs) to execute computation-intensive AI tasks (e.g., pedestrian detection for autonomous vehicles [1, 2]) through task-oriented computation offloading. This paradigm offloads computations from EDs to proximal edge servers (ESs) to optimize task-specific performance metrics like detection accuracy. However, transmitting high-fidelity raw data over bandwidth-limited wireless links incurs significant latency. To address this, EDs may perform data degradation before transmission. As Fig. 1 illustrates for visual object detection, EDs can select from multiple degradation levels and downsample high-resolution (e.g., 4K) video streams to lower-resolution (e.g., 720p or 480p). While this degradation reduces transmission latency, it sacrifices inference accuracy, creating an inherent accuracy-latency tradeoff. To achieve an optimal tradeoff under time-varying channel and content, an adaptive offloading scheme is required.
@@ II Related Works
Computation offloading is a well-established technique in mobile edge computing (MEC), enabling resource-constrained EDs to leverage a nearby ES for executing demanding tasks. Conventional MEC offloading strategies primarily follow two paradigms: binary offloading [9, 10, 11], where tasks are executed entirely locally or remotely, and partial offloading [12, 13, 14, 15, 16], which partitions computational workloads between EDs and the ES. Despite effectively optimizing system-level metrics (e.g., computation rate, energy consumption) using explicit objectives, these data-centric approaches treat task data as generic payloads, ignoring its critical role as content carrier for AI performance. As a result, they disregard how processing operations (e.g, resolution degradation) alter input content and consequently degrade task-specific performance such as inference accuracy. Conventional offloading strategies [9, 10, 11, 12, 13, 14, 15, 16] therefore fail to optimize the accuracy-latency tradeoff in EI systems.
@@ III-B Problem Formulation
During real-time inference, the detection accuracy $c_{t,n}$ remains unavailable due to the absence of ground truth $\mathcal{B}^{\rm{GT}}_{t,n}$. To address this limitation, we leverage the confidence values associated with predicted bounding boxes $\hat{\mathcal{B}}_{t,n}$ from the object detection model. The cumulative confidence metric $\alpha_{t,n}$, defined as the sum of confidence values across all predicted bounding boxes $\hat{\mathcal{B}}_{t,n}$, provides a real-time accuracy proxy strongly correlated with actual detection performance [22]. This enables us to reformulate the following alternative optimization problem:
@@ IV-A1 Probabilistic GP Modeling
The effectiveness of the above GP model critically depends on the kernel function $\kappa(\cdot,\cdot)$, which encodes prior information about the structure of utility function $U(\bm{z_{t}})$. To capture the complex interactions across the heterogeneous components of $\bm{z}_{t}=(\bm{h}_{t},\bm{a}_{t},t)$, we design a composite kernel that strategically combines three specialized kernels: a radial basis function (RBF) kernel $\kappa_{\bm{\theta}_{\rm h}}^{\rm{rbf}}$ with parameter $\bm{\theta}_{\rm h}$ for continuous channel states $\bm{h}_{t}$, a categorical kernel $\kappa_{\upsilon_{\rm a}}^{\rm{cat}}$ with parameter $\upsilon_{\rm a}$ for discrete degradation actions $\bm{a}_{t}$, and a temporal kernel $\kappa_{\rho}^{\rm{tmp}}$ with parameter $\rho$ for temporal dynamics. The complete mathematical definitions of these kernels and their parameters are provided in Appendix A. The resulting kernel structure of $\kappa(\cdot,\cdot)$ is formulated as:
@@
where $i,i^{\prime}\in\{t-J_{\rm B},\cdots,t-1\}$ are time indices.
@@ IV-B Customized Optimal Bandwidth Allocation under Fixed Degradation Action
$a_{t,n}\in\mathcal{A},b_{t,n}\geq 0,\tau_{t,n}^{\rm o}\geq 0,\forall n,$ (27d)
where $\bm{\tau}_{t}^{\rm o}=\{\tau_{t,n}^{\rm o},\forall n\}$. While standard convex optimization methods (e.g., interior point method [26]) efficiently solve (27), we derive a semi-closed-form solution using Lagrangian duality to characterize optimal resource allocation. Let $\bm{\phi}=\{\phi_{n},\forall n\}\geq 0$ and $\eta\geq 0$ denote Lagrangian multipliers for constraints (27b) and (27c), respectively. The partial augmented Lagrangian of (27) is
@@
Please see Appedix B for detail. ∎
@@ VII Simulation Results
To model autonomous driving scenarios, ED mobility follows a 100 m $\times$ 50 m rectangular trajectory centered at the ES location. The system operates over $T=3,000$ discrete time slots. At $t=0$, initial positions of EDs are randomly distributed along a designated 100-meter edge of the rectangle. Between consecutive time slots, ED positions progress in discrete 2.5-meter increments counterclockwise along the rectangular trajectory. We adopt Rayleigh fading channels for wireless communications between EDs and ES. The channel gain for EDn at time slot $t$ is modeled as $h_{t,n}=\varsigma\cdot\bar{h}_{t,n}$, where $\varsigma$ follows an exponential distribution with unit mean. The average channel gain $\bar{h}_{t,n}$ is defined as $\bar{h}_{t,n}=G_{A}\left(\frac{3\times 10^{8}}{4\pi f_{\rm c}d_{t,n}}\right)^{\lambda}$, where $G_{A}=4.11$ denotes antenna gain, $f_{\rm c}=2.4$ GHz is the carrier frequency, $d_{t,n}$ represents the Euclidean distance between EDn and ES at time $t$, and $\lambda=2.4$ by default is the path-loss exponent. We consider homogeneous EDs with identical transmit power $p_{n}=0.1$ watt, and weighting factors $w_{n}=w=1$ by default, $\forall n$. The total bandwidth is $W=5$ MHz. The noise power spectrum density at the ES is $\delta^{2}=-174$ dBm/Hz.
For the proposed framework LAB, the actor module employs a Transformer-based DNN consisting of three core components: a fully-connected input embedding layer that generates 64-dimensional embeddings, a 2-layer Transformer encoder with 4 attention heads per layer and 256-dimensional feed-forward networks, and a fully-connected output layer using sigmoid activation to produce $\hat{\bm{\Omega}}_{t}$. The default parameters of DNN include: input state history length $l=1$, replay memory size $J_{\rm D}=512$, minibatch size $J_{\rm D}^{\rm s}=128$, initial number of candidates $K^{0}=\min\{8N,A^{N}\}$ with update interval $\Delta_{\rm K}=32$, DNN training interval $\Delta_{\rm D}=20$, and the Adam optimizer learning rate $\xi=0.01$. The critic module employs the L-BFGS-B method to update the BO parameters, with defaults including observation cache size $J_{\rm B}=256$ and BO update interval $\Delta_{B}=20$. All the simulations are conducted on an Intel(R) Xeon(R) Silver 4110 (2.10 GHz) machine with a Tesla P40 GPU. The complete source code for LAB will be published on GitHub upon acceptance.
@@ VII-A Parameter Sensitivity Analysis of LAB
In Fig. 4, we demonstrate the confidence-latency trade-off of LAB under different delay weighting factors. Specifically, we compute the long-term average confidence $\bar{\alpha}=\frac{1}{T}\sum_{t=0}^{T-1}\sum_{n=1}^{N}\alpha_{t,n}$ and the long-term average latency $\bar{\tau}=\frac{1}{T}\sum_{t=0}^{T-1}\sum_{n=1}^{N}\tau_{t,n}$. As $w$ increases from 0 to $3$, LAB increasingly prioritizes latency reduction, resulting in lower $\bar{\tau}$ at the expense of reduced $\bar{\alpha}$. Notably, increasing $w$ from 0 to 1 reduces $\bar{\tau}$ dramatically by 68.3% (from 1.83 s to 0.58 s) while only decreasing $\bar{\alpha}$ by 6.9% (from 5.21 to 4.85). Therefore, we select $w=1$ for all subsequent simulations as it maintains high confidence without incurring substantial latency penalties.
@@
Figure 5: Optimality gap under different LAB parameters: (a) acqusition function; (b) BO cache size; (c) history length of input sate; and (d) update interval.
@@
In Fig. 5, we analyze parameter sensitivity of LAB through the optimality gap metric $U_{t}^{\ast}-U_{t}$, where $U_{t}^{\ast}$ represents the theoretical maximum utility obtained through exhaustive search over all $A^{N}$ degradation actions for (15), with results window-averaged over 500 time slots. Fig. 5(a) reveals that UCB acquisition function consistently outperforms expected improvement (EI) and probability of improvement (PI) alternatives by more effectively balancing exploration-exploitation trade-offs. For the BO cache in Fig. 5(b), a size of $J_{\rm B}=256$ achieves the minimal optimality gap, while a smaller cache ($J_{\rm B}=128$) exhibits degraded performance due to insufficient historical data utilization, and a larger cache ($J_{\rm B}=512$) introduces noisy temporal correlations that compromise decision quality. State history configuration in Fig. 4(c) reveals that $l=1$ (current state only) delivers optimal performance, whereas extended histories (e.g., $l=10$) provides only marginal asymptotic improvement at increased computational overhead. Besides, we notice from Fig. 4(d) that while reducing $\Delta_{\rm D}$ and $\Delta_{\rm B}$ generally lowers optimality gaps, the improvement plateaus below 20 time slots. Further reduction to $\Delta_{\rm D}=\Delta_{\rm B}=10$ provides only marginal gains while substantially increasing computational load. To balance system performance with implementation efficiency, we adopt UCB acquisition, $J_{\rm B}=256$, $l=1$, and $\Delta_{\rm D}=\Delta_{\rm B}=20$ as the default configuration.
VII-B System Performance Comparison

### v2

@@ Abstract
Edge intelligence (EI) allows resource-constrained edge devices (EDs) to offload computation-intensive AI tasks (e.g., visual object detection) to edge servers (ESs) for fast execution. However, transmitting high-volume raw task data (e.g., 4K video) over bandwidth-limited wireless networks incurs significant latency. While EDs can reduce transmission latency by degrading data before transmission (e.g., reducing resolution from 4K to 720p or 480p), it often deteriorates inference accuracy, creating a critical accuracy-latency tradeoff. The difficulty in balancing this tradeoff stems from the absence of closed-form models capturing content-dependent accuracy-latency relationships. Besides, under bandwidth sharing constraints, the discrete degradation decisions among the EDs demonstrate inherent combinatorial complexity. Mathematically, it requires solving a challenging black-box mixed-integer nonlinear programming (MINLP). To address this problem, we propose LAB, a novel learning framework that seamlessly integrates deep reinforcement learning (DRL) and Bayesian optimization (BO). Specifically, LAB employs: (a) a DNN-based actor that maps input system state to degradation actions, directly addressing the combinatorial complexity of the MINLP; and (b) a BO-based critic with an explicit model built from fitting a Gaussian process surrogate with historical observations, enabling model-based evaluation of degradation actions. For each selected action, optimal bandwidth allocation is then efficiently derived via convex optimization. Numerical evaluations on real-world self-driving datasets demonstrate that LAB achieves near-optimal accuracy-latency tradeoff, exhibiting only 1.22% accuracy degradation and 0.07s added latency compared to exhaustive search. Notably, it outperforms conventional DRL with 3.29% higher accuracy and 42.60% lower latency, demonstrating its advantageous performance in handling black-box optimization problems. The complete source code for LAB is available at https://github.com/Ethan-Xian-Li/LAB.
@@ I Introduction
Edge inference (EI) enables resource-constrained edge devices (EDs) to execute computation-intensive AI tasks (e.g., pedestrian detection and cooperative perception for autonomous vehicles [2, 3, 4]) through task-oriented computation offloading. This paradigm offloads computations from EDs to proximal edge servers (ESs) to optimize task-specific performance metrics like detection accuracy. However, transmitting high-fidelity raw data over bandwidth-limited wireless links incurs significant latency. To address this, EDs may perform data degradation before transmission. As Fig. 1 illustrates for visual object detection, EDs can select from multiple degradation levels and downsample high-resolution (e.g., 4K) video streams to lower-resolution (e.g., 720p or 480p). While this degradation reduces transmission latency, it sacrifices inference accuracy, creating an inherent accuracy-latency tradeoff. To achieve an optimal tradeoff under time-varying channel and content, an adaptive offloading scheme is required.
@@ II Related Works
Computation offloading is a well-established technique in mobile edge computing (MEC), enabling resource-constrained EDs to leverage a nearby ES for executing demanding tasks. Conventional MEC offloading strategies primarily follow two paradigms: binary offloading [11, 12, 13, 14], where tasks are executed entirely locally or remotely, and partial offloading [15, 16, 17, 18, 19], which partitions computational workloads between EDs and the ES. Despite effectively optimizing system-level metrics (e.g., computation rate, energy consumption) using explicit objectives, these data-centric approaches treat task data as generic payloads, ignoring its critical role as content carrier for AI performance. As a result, they disregard how processing operations (e.g, resolution degradation) alter input content and consequently degrade task-specific performance such as inference accuracy. Conventional offloading strategies [11, 12, 13, 14, 15, 16, 19, 17, 18] therefore fail to optimize the accuracy-latency tradeoff in EI systems.
@@ III-B Problem Formulation
During real-time inference, the detection accuracy $c_{t,n}$ remains unavailable due to the absence of ground truth $\mathcal{B}^{\rm{GT}}_{t,n}$. To address this limitation, we leverage the confidence values associated with predicted bounding boxes $\hat{\mathcal{B}}_{t,n}$ from the object detection model. The cumulative confidence metric $\alpha_{t,n}$, defined as the sum of confidence values across all predicted bounding boxes $\hat{\mathcal{B}}_{t,n}$, provides a real-time accuracy proxy strongly correlated with actual detection performance [25]. This enables us to reformulate the following alternative optimization problem  [^1] While it admits an equivalent per-slot decomposition, the long-term formulation in (13) represents a general starting point. It provides a consistent framework that can naturally accommodate future extensions involving inter-slot dependencies, e.g., device energy budgets and queue stability.:
@@ IV-A1 Probabilistic GP Modeling
The effectiveness of the above GP model critically depends on the kernel function $\kappa(\cdot,\cdot)$, which encodes prior information about the structure of utility function $U(\bm{z_{t}})$. To capture the complex interactions across the heterogeneous components of $\bm{z}_{t}=(\bm{h}_{t},\bm{a}_{t},t)$, we design the following composite kernel
@@
where $i,i^{\prime}\in\{t-J_{\rm B},\cdots,t-1\}$ are time indices. The constituent kernels $\kappa_{\bm{\theta}_{\rm h}}^{\rm{rbf}}$, $\kappa_{\upsilon_{\rm a}}^{\rm{cat}}$, and $\kappa_{\rho}^{\rm{tmp}}$ are defined as follows:
  • Channel State Kernel $\kappa_{\bm{\theta}_{\rm h}}^{\rm{rbf}}$: To capture the similarity in continuous channel states $\bm{h}_{t}$, we utilize an automatic relevance determination (ARD) radial basis function (RBF) kernel:
  • Degradation Action Kernel $\kappa_{\upsilon_{\rm a}}^{\rm{cat}}$,: To measure the similarity between discrete degradation actions $\bm{a}_{t}$, we implement a categorical kernel [30]:
  • Temporal Dynamics Kernel $\kappa_{\rho}^{\rm{tmp}}$: To model time-varying correlations arising from evolving channel conditions and content characteristics, we employ a temporal kernel [30]:
@@ IV-B Customized Optimal Bandwidth Allocation under Fixed Degradation Action
$b_{t,n}\geq 0,\tau_{t,n}^{\rm o}\geq 0,\forall n,$ (30d)
where $\bm{\tau}_{t}^{\rm o}=\{\tau_{t,n}^{\rm o},\forall n\}$. While standard convex optimization methods (e.g., interior point method [29]) efficiently solve (30), we derive a semi-closed-form solution using Lagrangian duality to characterize optimal resource allocation. Let $\bm{\phi}=\{\phi_{n}\geq 0,\forall n\}$ and $\eta\geq 0$ denote Lagrangian multipliers for constraints (30b) and (30c), respectively. The partial augmented Lagrangian of (30) is
@@
[…]
By setting (36) to 0, we obtain
$\tau_{t,n}^{\rm o\ast}=\sqrt{\frac{\phi_{n}^{\ast}d_{t,n}}{w_{n}}}.$ (38)
By setting (37) to 0, we have
By solving (39), we obtain that
where $\mathcal{W}(\cdot)$ is the Lambert-$\mathcal{W}$ function. ∎
@@ VII Simulation Results
To model autonomous driving scenarios, ED mobility follows a 100 m $\times$ 50 m rectangular trajectory centered at the ES location. The system operates over $T=3,000$ discrete time slots. At $t=0$, initial positions of EDs are randomly distributed along a designated 100-meter edge of the rectangle. Between consecutive time slots, ED positions progress in discrete 2.5-meter increments counterclockwise along the rectangular trajectory. We adopt Rayleigh fading channels for wireless communications between EDs and ES. The channel gain for EDn at time slot $t$ is modeled as $h_{t,n}=\varsigma\cdot\bar{h}_{t,n}$, where $\varsigma$ follows an exponential distribution with unit mean. The average channel gain $\bar{h}_{t,n}$ is defined as $\bar{h}_{t,n}=G_{A}\left(\frac{3\times 10^{8}}{4\pi f_{\rm c}d_{t,n}}\right)^{\lambda}$, where $G_{A}=4.11$ denotes antenna gain, $f_{\rm c}=2.4$ GHz is the carrier frequency, $d_{t,n}$ represents the Euclidean distance between EDn and ES at time $t$, and $\lambda=2.4$ by default is the path-loss exponent. We consider homogeneous EDs with identical transmit power $p_{n}=0.1$ watt, and weighting factors $w_{n}=w=1$ by default, $\forall n$. The total bandwidth is $W=5$ MHz. The noise power spectrum density at the ES is $\delta^{2}=-174$ dBm/Hz. In practice, the degradation latency $\tau_{t,n}^{d}$ in (2) and the edge computation latency $\tau_{t,n}^{c}$ in (9) are hard to model precisely with closed-form functions under dynamic hardware conditions. Instead, we measure these latencies in real time during execution.
For the proposed framework LAB, the actor module employs a Transformer-based DNN consisting of three core components: a fully-connected input embedding layer that generates 64-dimensional embeddings, a 2-layer Transformer encoder with 4 attention heads per layer and 256-dimensional feed-forward networks, and a fully-connected output layer using sigmoid activation to produce $\hat{\bm{\Omega}}_{t}$. The default parameters of DNN include: input state history length $l=1$, replay memory size $J_{\rm D}=512$, minibatch size $J_{\rm D}^{\rm s}=128$, initial number of candidates $K^{0}=\min\{8N,A^{N}\}$ with update interval $\Delta_{\rm K}=32$, DNN training interval $\Delta_{\rm D}=20$, and the Adam optimizer learning rate $\xi=0.01$. The critic module employs the L-BFGS-B method to update the BO parameters, with defaults including observation cache size $J_{\rm B}=256$ and BO update interval $\Delta_{B}=20$. The exploration coefficient in the UCB acquisition function (27) is $\zeta=\sqrt{0.2}$. Remarkably, even with this non-optimized $\zeta$ setting, our proposed LAB framework consistently outperforms both the standalone BO and state-of-the-art (SOTA) DRL baselines, as demonstrated later in Sec VII-C. All the simulations are conducted on an Intel(R) Xeon(R) Silver 4110 (2.10 GHz) machine with a Tesla P40 GPU.
@@ VII-A Parameter Sensitivity Analysis of LAB
In Fig. 4, we demonstrate the confidence-latency trade-off of LAB under different delay weighting factors. Specifically, we compute the long-term average confidence $\bar{\alpha}=\frac{1}{T}\sum_{t=0}^{T-1}\sum_{n=1}^{N}\alpha_{t,n}$ and the long-term average latency $\bar{\tau}=\frac{1}{T}\sum_{t=0}^{T-1}\sum_{n=1}^{N}\tau_{t,n}$. Note that a higher $\bar{\alpha}$ generally indicates more or more confident detections. As $w$ increases from 0 to $3$, LAB increasingly prioritizes latency reduction, resulting in lower $\bar{\tau}$ at the expense of reduced $\bar{\alpha}$. Notably, increasing $w$ from 0 to 1 reduces $\bar{\tau}$ dramatically by 68.3% (from 1.83 s to 0.58 s) while only decreasing $\bar{\alpha}$ by 6.9% (from 5.21 to 4.85). Therefore, we select $w=1$ for all subsequent simulations as it maintains high confidence without incurring substantial latency penalties.
@@
Figure 5: Optimality gap under different LAB parameters: (a) acqusition function; (b) BO cache size; (c) history length of input state; and (d) update interval.
@@
In Fig. 5, we analyze parameter sensitivity of LAB through the optimality gap metric $U_{t}^{\ast}-U_{t}$, where $U_{t}^{\ast}$ represents the theoretical maximum utility obtained through exhaustive search over all $A^{N}$ degradation actions for (15), with results window-averaged over 500 time slots. Fig. 5(a) reveals that UCB acquisition function consistently outperforms expected improvement (EI) and probability of improvement (PI) alternatives by more effectively balancing exploration-exploitation trade-offs. Note that the fluctuations in the optimality gap, such as the noticeable rise around $t=2000$, stems from the non-stationary nature of the real-world dataset. For the BO cache in Fig. 5(b), a size of $J_{\rm B}=256$ achieves the minimal optimality gap, while a smaller cache ($J_{\rm B}=128$) exhibits degraded performance due to insufficient historical data utilization, and a larger cache ($J_{\rm B}=512$) introduces noisy temporal correlations that compromise decision quality. State history configuration in Fig. 4(c) reveals that $l=1$ (current state only) delivers optimal performance, whereas extended histories (e.g., $l=10$) provides only marginal asymptotic improvement at increased computational overhead. Besides, we notice from Fig. 4(d) that while reducing $\Delta_{\rm D}$ and $\Delta_{\rm B}$ generally lowers optimality gaps, the improvement plateaus below 20 time slots. Further reduction to $\Delta_{\rm D}=\Delta_{\rm B}=10$ provides only marginal gains while substantially increasing computational load. To balance system performance with implementation efficiency, we adopt UCB acquisition, $J_{\rm B}=256$, $l=1$, and $\Delta_{\rm D}=\Delta_{\rm B}=20$ as the default configuration.
VII-B Ablation Study
Figure 6: Normalized long‑term average utility of LAB and its ablated variants: MAT ($\mathrm{Mat\acute{e}rn}$ kernel), OT (no temporal kernel), MLP (MLP-based actor), and X (BO with contextual input).
To validate the design choices of our LAB framework, we conduct an ablation study, with results presented in Fig. 6. We compare the long-term average utility of our default LAB against four ablated variants: LAB‑MAT (replaces the RBF kernel with a $\mathrm{Mat\acute{e}rn}$ kernel), LAB‑OT (removes the temporal kernel), MLP‑LAB (substitutes the Transformer-based actor with a multi‑layer perceptron), and LAB‑X (augments the kernel with image-content features). The results are normalized to that of LAB. We observe utility drops of 1.64% for LAB-MAT, 2.45% for LAB-OT, and 6.26% for MLP-LAB, respectively. These results confirm the effectiveness of our kernel design and actor architecture. Notably, LAB-X exhibits a more pronounced degradation of 10.09%. This counter-intuitive result likely stems from the dynamic and non‑stationary image content in the real‑world dataset, where rapidly changing scenes make historical content information misleading rather than informative.
VII-C System Performance Comparison